OpenStudy (anonymous):

Sally worked the following equation on her paper. She checked her answer and found that is was incorrect. Identify her mistakes and explain to her how to correct them. (hint: There are 2 mistakes.)

OpenStudy (anonymous):

3(8x-2) = 46 - 2(12x+1) 24x-6 = 46 - 24x + 2 24x -6 = 48 - 24x 48-6 = 48 48x = 42 x=42/48 or 7/8

OpenStudy (anonymous):

24x-6 = 46 - 24x + 2 48-6 = 48

OpenStudy (hba):

oj

OpenStudy (anonymous):

First mistake: -2(12x+1)=-24x+2 It should be -24x-2

OpenStudy (anonymous):

Second mistake: 48-6=48

OpenStudy (anonymous):

because it wouldnt equal the same amount correct?

OpenStudy (anonymous):

can u help me solve this please Solve the following equation and show your work. 2x + 2(x + 3) = 2(x + 7)